# GRACE: Generating Socially Appropriate Robot Actions Leveraging LLMs and Human Explanations

### v1

@@ V-B3 Action Appropriateness using Explanations
Quantitative Results: We include baselines that follow two different approaches for generating socially appropriate robot actions, as detailed in Section V-A3. First, the left side of Table III shows the baselines that generated social appropriateness scores using the scene features. The results were obtained using all the MannersDB+ data (except one scene with insufficient annotations). These results show that GRACE outperforms the former methods by producing closer score values to the human ground truth, i.e., lower RMSE (STD values were 0.01 both for MLP [1] and GRACE), and it captures the trend of human scores the best by providing the highest PCC and CCC values. The second set of baselines includes generative models that correct LLM predictions without leveraging human explanations to allow us to assess the impact of including explanations in the prediction process. These results are reported on the right side of Table III for MannersDB+ by discarding samples that did not include explanations belonging to the explanation categories described in Section III-B. The results show that the proposed GRACE system performs the best based on average RMSE (STD values were 0.01 for all models), PCC, and CCC, showing the significance of human explanations while generating socially appropriate robot actions. An extension of GRACE with a noising mask did not change the results noticeably.
@@ VI Discussion and Conclusion
In this work, we propose GRACE, a full system to generate socially appropriate robot actions leveraging LLMs and human explanations. GRACE first classifies the scenes based on their certainty, and for ‘certain’ cases, it directly outputs LLM predictions. To handle ‘uncertain’ scenes, GRACE improves the LLM predictions with a conditional autoencoder network structure, which includes human explanations in the learning process. Our evaluations showed that GRACE successfully learns the relationship between human scores and their explanations by outperforming several baselines and sucessfully generating sensible robot explanations.
@@
In addition to generating socially appropriate robot actions, GRACE is also capable of providing explanations for human scores as demonstrated in Figure 3. The generated explanations include fundamental scene attributes, such as crowd size, the robot’s proximity to others, human states, and the safety of actions, despite the system not being explicitly trained on scene features. These capabilities are crucial for HRI, enabling robots to provide reasoning in uncertain situations, boosting trust [28, 29], and improving human-robot collaboration [36].
With the uncertainty classification component, GRACE addresses the challenge of predicting inherent dataset uncertainty, where annotators gave differing scores for the same image. This task is difficult even for humans, as shown by the baseline chance-level accuracies in Table I. In parallel, the models also struggled with the task on both datasets. Still, MannersDB results outperformed MannersDB+, likely due to its higher average annotators per scene (14.73 vs. 3.08), resulting in more reliable variance. This suggests that increasing the number of annotations may improve uncertainty classification by better capturing human score agreement.
Future work can extend GRACE in several ways: i) The uncertainty classification performances can be improved using probabilistic models (i.e., Bayesian Neural Networks [37] & Deep Ensemble Methods [38]); and ii) GRACE can be deployed on a robot to assess the effectiveness and accuracy of real-time interactions, which is the focus of our future work.
Acknowledgements
Funding: This work has been supported in part by Google under the GIG funding scheme. F. I. Dogan and H. Gunes have been supported by EPSRC/UKRI under grant ref. EP/R030782/1 (ARoEQ). U. Ozyurt and G. Cinar have been supported by the Erasmus Traineeship program.
Open Access: For open access purposes, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
Data access: This study involves secondary analyses of the existing datasets, that are described and cited in the text.

### v2

@@ V-B3 Action Appropriateness using Explanations
Quantitative Results: We include baselines that follow two different approaches for generating socially appropriate robot actions, as detailed in Section V-A3. First, the left side of Table III shows the baselines that generated social appropriateness scores using the scene features. The results were obtained using all the MannersDB+ data (except one scene with insufficient annotations). These results show that GRACE outperforms the former methods by producing closer score values to the human ground truth, i.e., lower RMSE (STD: 0.01 both for MLP [1] and GRACE), and it captures the trend of human scores the best by providing the highest PCC and CCC values. The second set of baselines includes generative models that correct LLM predictions without leveraging human explanations to allow us to assess the impact of including explanations in the prediction process. These results are reported on the right side of Table III for MannersDB+ by discarding samples that did not include explanations belonging to the explanation categories described in Section III-B. The results show that the proposed GRACE system performs the best based on average RMSE (STD: 0.01 for all models), PCC, and CCC, showing the significance of human explanations while generating socially appropriate robot actions. An extension of GRACE with a noising mask did not change the results noticeably.
@@ VI Discussion and Conclusion
In this work, we propose GRACE, a full system to generate socially appropriate robot actions leveraging LLMs and human explanations. GRACE first classifies the scenes based on their certainty, and for ‘certain’ cases, it directly outputs LLM predictions. To handle ‘uncertain’ scenes, GRACE improves the LLM predictions with a conditional autoencoder network structure, which includes human explanations in the learning process. Our evaluations showed that GRACE successfully learns the relationship between human scores and their explanations by outperforming several baselines and successfully generating sensible robot explanations.
@@
In addition to generating socially appropriate robot actions, GRACE can provide sensible explanations for human-provided scores (Figure 3). These explanations include critical scene attributes, such as crowd size, the robot’s proximity to others, human states, and the safety of actions, despite the system not being explicitly trained on scene features. These capabilities are crucial for HRI, enabling robots to provide reasoning in uncertain situations, boosting trust [28, 29], and improving human-robot collaboration [36].
With the uncertainty classification, GRACE addresses the challenge of predicting inherent dataset uncertainty, where annotators gave differing scores for the same image. This task is difficult even for humans and similarly challenging for ML models, as shown in Table I. Such misclassifications could potentially cause GRACE to directly output LLM predictions for uncertain cases or ask for explanations even in certain situations. Although the models struggled with the task, MannersDB results outperformed MannersDB+, likely due to its higher average annotators per scene (14.73 vs. 3.08), resulting in more reliable variance. Hence, increasing the number of annotations may improve uncertainty classification by better capturing human score agreement.
Future work can extend GRACE in several ways: i) The uncertainty classification performances can be improved using probabilistic models (i.e., Bayesian Neural Networks [37] & Deep Ensemble Methods [38]); and ii) GRACE can be deployed on a robot to assess the effectiveness and accuracy of real-time interactions, i.e., the focus of our future work.